OpenStudy (candyme):

Solve for n and show your work. 2/3n+6=1/4n-3

OpenStudy (anonymous):

Regroup and reduce.

OpenStudy (candyme):

I know how to solve this type, but I have done it several times and gotten it wrong.

OpenStudy (anonymous):

So, try to get it so that n is only on one side of the equation. Any idea how we can do that?

OpenStudy (anonymous):

Okay, well walk me through your steps. We'll figure out where you've gone wrong.

OpenStudy (candyme):

First I made the two fractions the same denominator so I could group them. It turned into 8/12n+6=3/12n-3. +3/12 +3/12 11/12n+6=-3 -6 -6 11/12n=-9 n=-9.81818181

OpenStudy (anonymous):

You changed the denominator correctly, but check your second step.

OpenStudy (anonymous):

8/12 + 3/12 = 11/12 like you said. However, 3/12 + 3/12 =0?

OpenStudy (candyme):

OH now I see... I should have subtracted it so it would be 5/12n+6=-3 -6 -6 -9 5/12n=-9 n=-21.6 Thanks!

OpenStudy (anonymous):

My pleasure =) Good work.

OpenStudy (anonymous):

No he was right. Multiply both sides by 12/5. 5/12* n * 12/5 = -9 * (12/5) (12*5)/(12*5)*n = -108/5 1*n = -21.6 n = -21.6 I think the n was never meant to be in the denominator.

OpenStudy (candyme):

I got it right with -21.6 Thanks to all who helped :)

OpenStudy (candyme):

No,the equation is 2/3 n +6= 1/4 n -3